# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

### v1

@@ 2. Preliminaries
Note that ${\rm Re\ }\bigg{(}\frac{p_{M}(\xi)}{e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}\bigg{)}>0$ for $|\xi|^{\frac{1}{3}}M>0$ by numberical results (see figure 2, where $r=|\xi|^{\frac{1}{3}}M$). Then by continuity of $m(\xi)$, we have $\frac{1}{|m(\xi)|}\leq C(1+|\xi|)^{-\frac{4}{3}}$ for any $\xi\in{\mathbb{R}}$ and $M>0$, where $C>0$ depends on $M^{-1}$.
Figure 2. Analysis about $m(\xi)$.
4. Proof of Main Theorem
Let us recall the perturbation system (1.15) that
(4.1) $\left\{\begin{aligned} &y\omega_{x}-\omega_{yy}=-u\omega_{x}-v\omega_{y},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\\ &\omega_{y}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x).\end{aligned}\right$.
In order to find a solution of the equation (4.1), we use the iteration method. Set
@@ 4. Proof of Main Theorem
Then, we reduce (4.1) to the following BVP:
@@
Thus, we can consider the iterated system as following:
@@
Recall that the space $X_{\alpha,\infty}$ is defined as
(4.4) $X_{\alpha,\infty}=\bigg{\{}f:{\mathbb{R}}\to{\mathbb{R}}\big{|}\ \ {\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}f\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big{\|}|\partial_{x}|^{\frac{5}{6}}f\big{\|}}_{L^{2}({\mathbb{R}})}+{\big{\|}f\big{\|}}_{L^{\infty}({\mathbb{R}})}<\infty\bigg{\}}.$
First of all, we give some results about the iterated system (4.3).
@@
Given $\alpha\in(\frac{2}{3},\frac{7}{3}]$. Suppose $k$ is a non-negative integer and $(\bar{\omega}_{k},\bar{A}_{k})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Then there exists a solution $(\bar{\omega}_{k+1},\bar{A}_{k+1})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ of (4.3). Moreover, for $\varepsilon\in(0,\frac{1}{6}]$, we have
@@
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}\big{[}\big{(}{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}\big{)}^{2}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\big{]}.$
@@
Set $\bar{h}_{k}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0})$, we decompose $\bar{\omega}_{k+1}=\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}$, where $\bar{\omega}_{e,k+1}$ and $\bar{\omega}_{b,k+1}$ satisfy
(4.6) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e,k+1}-\partial_{y}^{2}\bar{\omega}_{e,k+1}=\bar{h}_{k},\\ &\partial_{y}\bar{\omega}_{e,k+1}\big{|}_{y=0}=0,\quad\bar{\omega}_{e,k+1}\big{|}_{y=M}=0,\end{aligned}\right$.
$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b,k+1}-\partial_{y}^{2}\bar{\omega}_{b,k+1}=0,\\ &\partial_{y}\bar{\omega}_{b,k+1}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1}.\end{aligned}\right$.
with
@@
By Lemma 3.1, we obtain the unique solution $\bar{\omega}_{e,k+1}$ with the following form
@@
By Lemma 3.2, we can find a solution $\bar{\omega}_{b,k+1}$ as following:
(4.9) $\hat{\omega}_{b,k+1}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0.\end{aligned}\right$.
@@
By Lemma 3.1, Theorem 1.1, and Proposition 2.3, we obtain
@@
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}({\big{\|}\bar{\omega}_{0}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}\partial_{x}A_{0}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}$
$\qquad+{\big{\|}\partial_{x}A_{k}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}+{\big{\|}\partial_{x}F\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2})$
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}({\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}A_{k}\big{\|}}_{X_{\alpha,\infty}}^{2}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}).$
By Lemma 3.2 and Proposition 2.4, we have
(4.12) ${\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{\alpha,\infty}}={\big{\|}|\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}$
[…]
$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}},$
${\big{\|}\bar{\omega}_{b,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}$
$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Combining (4.11) with (4.12), one gets
$\qquad\leq{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{\omega}_{b,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}\bigg{[}({\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}A_{k}\big{\|}}_{X_{\alpha,\infty}}^{2})+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\bigg{]}.$
Then it completes the proof.
To get the uniform estimates of the solutions, we use the following results which can be got the induction method.
@@
(4.14) $(1)\ a_{k+1}\leq\alpha_{1}a_{k}^{2}+\alpha_{2},\quad\mbox{for any }k\geq 0;\quad(2)\ b_{k+1}\leq\beta_{1}b_{k}+\beta_{2}b_{k}^{2},\quad\mbox{for any }k\geq 1.$
Assume that $\alpha_{2}\leq\frac{1}{4\alpha_{1}}$ and $\beta_{1}<1$. Set $a_{0}=0$ and $b_{1}\leq\frac{\gamma-\beta_{1}}{\beta_{2}}$ for some $\gamma\in(\beta_{1},1)$, then we have
(4.15) $0\leq\sup_{k\geq 0}a_{k}\leq 2\alpha_{2},\qquad\sum_{k=1}^{\infty}b_{k}\leq\frac{b_{1}}{1-\gamma}.$
Now, we are in the position to give the proof of Theorem 1.2.
@@
The proof proceeds in three steps.
Step 1. Bounded sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$.
Taking $\bar{\omega}_{0}=\bar{A}_{0}=0$, by Proposition 4.1, we obtain a sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Set $a_{k}={\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}$. Clearly, we have $a_{0}=0$. By Proposition 4.1, one gets
(4.16) $a_{k+1}\leq CM^{\frac{5+3\varepsilon}{2}}a_{k}^{2}+CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}.$
Using Lemma 4.1, for $4C^{2}M^{5+3\varepsilon}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\leq 1$ (that is, ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$), we obtain
(4.17) $\sup_{k\geq 0}\bigg{(}{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}=\sup_{k\geq 0}a_{k}\leq 2CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}.$
Step 2. Convergence of this sequence.
@@
which means ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$ for $M\gg 1$. Then there holds
@@
[…]
Step 3. Uniqueness of $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ in ${\mathcal{X}}_{\alpha,M}$.
Recall that
$\qquad\qquad\quad\omega(\cdot,y)\big{|}_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A\bigg{\}}.$
By Theorem 1.1, we obtain $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$. Combing (4.8)-(4.9) with the convergence of $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $X_{0}\cap X_{\alpha}\times X_{\alpha,\infty}$, we have $(\bar{\omega},\bar{A})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ and
(4.22) $\hat{\bar{\omega}}(\xi,y)\big{|}_{y=M}=\hat{\bar{\omega}}_{b}(\xi,y)\big{|}_{y=M}$
$\qquad=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0,\end{aligned}\right$.
which implies $(\bar{\omega},\bar{A})\in{\mathcal{X}}_{\alpha,M}$. Moreover, it is evident to verify that $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}-\{0\}$ due to $F\not\equiv 0$, which implies the space ${\mathcal{X}}_{\alpha,M}$ is non-trivial. Let $(\omega_{i},A_{i})\ (i=1,2)\in{\mathcal{X}}_{\alpha,M}$ be two solutions of the perturbation system (1.15). Set $\tilde{\omega}=\omega_{1}-\omega_{2},\ \tilde{A}=A_{1}-A_{2}$, then there holds
[…]
Similar to the proof of Proposition 4.1, we decompose $\tilde{\omega}=\tilde{\omega}_{e}+\tilde{\omega}_{b}$ with
(4.24) $\left\{\begin{aligned} &y\partial_{x}\tilde{\omega}_{e}-\partial_{y}^{2}\tilde{\omega}_{e}=\tilde{h},\\ &\partial_{y}\tilde{\omega}_{e}\big{|}_{y=0}=0,\quad\tilde{\omega}_{e}\big{|}_{y=M}=0,\end{aligned}\right$.
[…]
By Lemma 3.1, we have
@@
By Lemma 3.2, we obtain
(4.26) $\hat{\tilde{\omega}}_{b}(\xi,y)=C(\xi)\bigg{(}\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\bigg{)}+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
In view of $\hat{\tilde{\omega}}_{e}\big{|}_{y=M}=0$, one gets
(4.27) $\hat{\tilde{\omega}}_{b}\big{|}_{y=M}=\hat{\tilde{\omega}}\big{|}_{y=M}=\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)}(i\xi)^{\frac{2}{3}}|\xi|\hat{\tilde{A}}$.
Combining (4.26) with (4.27), we obtain
@@
(4.28) $\hat{\tilde{\omega}}_{b}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
This formula is identical to (3.36), then using the Lemma 3.2, we obtain
@@
Since $I_{M}[\tilde{\omega}_{e}+\tilde{\omega}_{b}]=\tilde{A}$, we have
(4.30) $\tilde{A}=m(-i\partial_{x})^{-1}I_{M}[\tilde{\omega}_{e}].$
It follows from Proposition 2.4 that

### v2

@@ 3. The solvability of the linear system
Note that ${\rm Re\ }\bigg(\frac{p_{M}(\xi)}{e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}\bigg)>0$ for $|\xi|^{\frac{1}{3}}M>0$ by numberical results (see figure 2, where $r=|\xi|^{\frac{1}{3}}M$). Then by the continuity of $m(\xi)$, we have
$\frac{1}{|m(\xi)|}\leq C(1+|\xi|)^{-\frac{4}{3}}\text{ for any }\xi\in{\mathbb{R}}\text{ and }M>0,$
where $C>0$ depends on $M^{-1}$.
Figure 2. Analysis on $m(\xi)$.
4. Existence and uniqueness of classical solution
In this section, we establish the existence and uniqueness of classical solution to the nonlinear system (1.13), i.e.,
(4.1) $\left\{\begin{aligned} &y\partial_{x}\omega-\partial_{y}^{2}\omega=-u\partial_{x}\omega-v\partial_{y}\omega,\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\\ &\partial_{y}\omega\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x).\end{aligned}\right$.
To construct a solution to the system (4.1), we use the iteration method. Let
@@ 4. Existence and uniqueness of classical solution
Then the system (4.1) is reduced to the following BVP:
@@
We introduce the iterated system as follows
@@
First of all, we present the following uniform estimate for the iterated system (4.3).
@@
Given $\alpha\in(\frac{2}{3},\frac{7}{3}]$, assume that $(\bar{\omega}_{k},\bar{A}_{k})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Then there exists a solution $(\bar{\omega}_{k+1},\bar{A}_{k+1})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ to (4.3), which holds that for $\varepsilon\in(0,\frac{1}{6}]$,
@@
$\leq CM^{\frac{5+3\varepsilon}{2}}\Big[\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}\big)^{2}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\Big].$
@@
We decompose $\bar{\omega}_{k+1}=\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}$, where $\bar{\omega}_{e,k+1}$ and $\bar{\omega}_{b,k+1}$ satisfy
$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e,k+1}-\partial_{y}^{2}\bar{\omega}_{e,k+1}=\bar{h}_{k},\\ &\partial_{y}\bar{\omega}_{e,k+1}\big|_{y=0}=0,\quad\bar{\omega}_{e,k+1}\big|_{y=M}=0,\end{aligned}\right$.
and
$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b,k+1}-\partial_{y}^{2}\bar{\omega}_{b,k+1}=0,\\ &\partial_{y}\bar{\omega}_{b,k+1}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1},\end{aligned}\right$.
together with
@@
Here
$\bar{h}_{k}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0}).$
By Proposition 3.1, we obtain the unique solution $\bar{\omega}_{e,k+1}$ with the form
@@
By Proposition 3.2, we can find a solution $\bar{\omega}_{b,k+1}$ with
(4.7) $\hat{\omega}_{b,k+1}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),&\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),&xi<0.\end{aligned}\right$.
@@
By Proposition 3.1, Theorem 1.1 and Proposition 2.1, we obtain
@@
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}\big({\big\|\bar{\omega}_{0}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|\partial_{x}A_{0}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}$
$\qquad+{\big\|\partial_{x}A_{k}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}+{\big\|\partial_{x}F\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}\big)$
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|A_{k}\big\|}_{X_{\alpha,\infty}}^{2}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\big).$
By Proposition 3.2, Lemma 2.3 and (4.8), we have
(4.10) $\begin{split}{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}&={\big\||\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\\ &\leq C{\big\||\partial_{x}|^{\frac{5}{6}}I_{M}[\bar{\omega}_{e,k+1}]\big\|}_{L^{2}({\mathbb{R}})}+C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}I[\bar{\omega}_{e,k+1}]\big\|}_{L^{2}({\mathbb{R}})}\\ &\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}},\\ {\big\|\bar{\omega}_{b,k+1}\big\|}_{X_{0}\cap X_{\alpha}}&\leq C{\big\||\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}+C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\\ &\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}},\end{split}$
from which and (4.9), we arrive at
${\big\|\bar{\omega}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}$
$\leq{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{\omega}_{b,k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}$
$\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}$
$\leq CM^{\frac{5+3\varepsilon}{2}}\Big[\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|A_{k}\big\|}_{X_{\alpha,\infty}}^{2}\big)+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\Big].$
The proof is completed. ∎
To obtain the uniform estimates of the solution sequence, we use the following basic lemma, which can be proved by the induction.
@@
$a_{k+1}\leq\alpha_{1}a_{k}^{2}+\alpha_{2}\quad\mbox{for any }k\geq 0;\quad b_{k+1}\leq\beta_{1}b_{k}+\beta_{2}b_{k}^{2}\quad\mbox{for any }k\geq 1.$
Assume that $\alpha_{2}\leq\frac{1}{4\alpha_{1}}$ and $\beta_{1}<1$. Set $a_{0}=0$ and $b_{1}\leq\frac{\gamma-\beta_{1}}{\beta_{2}}$ for some $\gamma\in(\beta_{1},1)$. Then we have
$0\leq\sup_{k\geq 0}a_{k}\leq 2\alpha_{2},\qquad\sum_{k=1}^{\infty}b_{k}\leq\frac{b_{1}}{1-\gamma}.$
With the above preparations, we complete the proof of Theorem 1.2.
@@
The proof consists of three steps.
Step 1. Boundedness of the solution sequence.
Taking $\bar{\omega}_{0}=\bar{A}_{0}=0$, by Proposition 4.1, we obtain a sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Set $a_{k}={\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}$. Clearly, we have $a_{0}=0$. By Proposition 4.1, we have
$a_{k+1}\leq CM^{\frac{5+3\varepsilon}{2}}a_{k}^{2}+CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
By Lemma 4.1, for $4C^{2}M^{5+3\varepsilon}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\leq 1$ (that is, ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$), we obtain
$\sup_{k\geq 0}\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}\big)=\sup_{k\geq 0}a_{k}\leq 2CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
Step 2. Convergence of the solution sequence.
@@
[…]
which means that ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$ for $M\gg 1$, then there holds
@@
Step 3. Uniqueness of the solution $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ in ${\mathcal{X}}_{\alpha,M}$.
By Theorem 1.1, we know $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$. Combing (4.6)-(4.7) with the convergence of $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $X_{0}\cap X_{\alpha}\times X_{\alpha,\infty}$, we conclude that $(\bar{\omega},\bar{A})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ and
$\hat{\bar{\omega}}(\xi,y)\big|_{y=M}$ $=\hat{\bar{\omega}}_{b}(\xi,y)\big|_{y=M}$
$=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0.\end{aligned}\right$.
This shows that $(\bar{\omega},\bar{A})\in{\mathcal{X}}_{\alpha,M}$. Moreover, it is easy to verify that $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}\setminus\{0\}$ due to $F\not\equiv 0$, which implies that the space ${\mathcal{X}}_{\alpha,M}$ is non-trivial.
Let $(\omega_{i},A_{i})\ (i=1,2)\in{\mathcal{X}}_{\alpha,M}$ be two solutions to the perturbation system (1.13). Set $\tilde{\omega}=\omega_{1}-\omega_{2},\ \tilde{A}=A_{1}-A_{2}$, which solve
@@
As in the proof of Proposition 4.1, we decompose $\tilde{\omega}=\tilde{\omega}_{e}+\tilde{\omega}_{b}$ with
$\left\{\begin{aligned} &y\partial_{x}\tilde{\omega}_{e}-\partial_{y}^{2}\tilde{\omega}_{e}=\tilde{h},\\ &\partial_{y}\tilde{\omega}_{e}\big|_{y=0}=0,\quad\tilde{\omega}_{e}\big|_{y=M}=0,\end{aligned}\right$.
and
@@
By Proposition 3.1, we have
@@
By Proposition 3.2, we get
(4.14) $\hat{\tilde{\omega}}_{b}(\xi,y)=C(\xi)\big(\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\big)+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
In view of $\hat{\tilde{\omega}}_{e}\big|_{y=M}=0$, we have
$\hat{\tilde{\omega}}_{b}\big|_{y=M}=\hat{\tilde{\omega}}\big|_{y=M}=\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)}(i\xi)^{\frac{2}{3}}|\xi|\hat{\tilde{A}}$,
which along with (4.14) gives
@@
$\hat{\tilde{\omega}}_{b}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
This formula is identical to (3.34). We then get by Proposition 3.2 that
@@
Thanks to $I_{M}[\tilde{\omega}_{e}+\tilde{\omega}_{b}]=\tilde{A}$, we have
$\tilde{A}=m(-i\partial_{x})^{-1}I_{M}[\tilde{\omega}_{e}].$
It follows from Lemma 2.3 that